OpenStudy (cherie_magee):

Find the next term of the given sequence. 1.31, 2.54, 3.77, ... 4.90 5.00 5.10

OpenStudy (ivettef365):

just add 1.23 to the number

OpenStudy (ivettef365):

1.31 + 1.23 = 2.54 2.54 + 1.23 = 3.77 3.77 + 1.23 = ??

OpenStudy (cherie_magee):

5.00

OpenStudy (ivettef365):

correct